1. A desktop computer is to be cooled by a fan whose flow rate is 0.30 m3/min. Determine the mass flow rate of air through the fan at an elevation of 3400 m where the air density is 0.7 kg/m3. Also, if the average velocity of air is not to exceed 95 m/min, determine the minimum diameter of the casing of the fan. Answers: 0.00350 kg/s, 0.0634 m.



2. Air can a nozzle steadily at 2.21 kg/m3 and 20 m/s and leaves at 0.762 kg/m3 and 150 m/s. If the inlet area of the nozzle is 60 cm2, determine (a) the mass flow rate through the nozzle, and (b) the exit area of the nozzle. Answers: a) 0.265 kg/s, (b) 23.2 an'



1
Expert's answer
2021-12-14T15:58:32-0500

1.

a)

"\dot m=\rho\dot V=3.5~\frac gs,"

b)

"\dot V=\pi d^2v,\implies d=\sqrt{\frac{\dot V}{\pi v}}=3.2~cm."


2.

a)

"\dot Q=v_1S_1D_1=0.27~\frac{kg}s,"

b)

"S_2=\frac{\dot Q}{v_2D_2}=23.2~cm^2."